Hospital records show that of patients suffering from a certain disease, 25% die
of it.
(i) What is the probability that of 6 randomly selected patients, 4 will recover?
[3 marks]
(ii) What is the most probable number of recoveries out of 6 randomly selected
patients?
[
1
Expert's answer
2021-09-02T23:15:29-0400

Let "X=" the number of recovered patients: "X\sim Bin(n, p)."

Given "n=6, q=0.25, p=1-q=1-0.25=0.75."

(i)


"P(X=4)=\dbinom{6}{4}(0.75)^4(0.25)^{6-4}"

"=0.296630859375"

(ii) The mode is an integer "M"  that satisfies


"(n+1)p-1\leq M<(n+1)p"

"(6+1)\cdot0.75-1\leq M<(6+1)\cdot0.75"

"4.25\leq M<5.25"

"M=5"

"P(X=5)=\dbinom{6}{5}(0.75)^5(0.25)^{6-5}"

"=0.35595703125"

The most probable number of recoveries out of 6 randomly selected patients is "5" patients.